OpenStudy (anonymous):

Please Help!! The definite integral of the absolute value of sin2x-cos4x with the limit from 0 to pi/4

OpenStudy (thomas5267):

\[ \cos(4x)=1-2\sin^2(2x)\\ \sin(2x)-1+2\sin^2(2x)=0 2\sin^2(2x)+\sin^2(2x)-1=0\\ \] Solve the quadratic in sin(x) and figure out where do you have to flip the graph.